OpenStudy (anonymous):

cos(theta)tan(theta) + sin ( theta) / tan ( theta) = 2 cos( theta) should i start out with cos ( theta) sinx/ cosx + sin (theta)

OpenStudy (akashdeepdeb):

Are you sure you wrote the question properly? Because cosθ + sinθ is not equal to 2 cosθ! :/

OpenStudy (anonymous):

The question stated that cos(θ) tan (θ) + sin ( θ) / tan (θ) = 2 cos ( θ)

OpenStudy (anonymous):

For some reason i got 2 sin ( θ)

OpenStudy (loser66):

Question: how about the question? Is it "prove" or "solve"?

OpenStudy (akashdeepdeb):

I think the question is wee-bit wrong!

OpenStudy (anonymous):

It said derive the given identities

OpenStudy (loser66):

\[cos \theta tan\theta + \frac{sin\theta}{tan\theta}= 2 cos\theta ~~~~or~~~\frac{cos\theta tan\theta+sin\theta}{tan\theta} = 2 cos\theta\]

OpenStudy (anonymous):

the second one

OpenStudy (loser66):

hehehe... redo please @AkashdeepDeb

OpenStudy (anonymous):

hahaha

OpenStudy (akashdeepdeb):

:P Forget this the question is worng! sinθ + cosθ can never in the history of math be 2cosθ Unless θ = 45'

OpenStudy (loser66):

nope, it's right.!! you see, from the second one, you break it out, you get the first one which you think you are wrong, hahahaha... and it goes to the right handside.

OpenStudy (akashdeepdeb):

We have to SOLVE this for θ ??? Or PROVE IT????

OpenStudy (loser66):

prove.

OpenStudy (loser66):

ok, let me do

OpenStudy (akashdeepdeb):

OMG it is the WHOLE goddamn thing divided by tanθ IS IT??????????

OpenStudy (loser66):

\[\frac{cos \theta tan\theta +sin\theta}{tan\theta} = cos \theta + \frac{sin\theta}{tan\theta}= cos \theta + cos \theta = 2 cos\theta\]hehahahaha

OpenStudy (akashdeepdeb):

O.O Use Parenthesis @youndi :D

OpenStudy (loser66):

suggestion: practice Latex

OpenStudy (anonymous):

I wil. Thanks alot both of you. I appreciate it :)

OpenStudy (akashdeepdeb):

:)

OpenStudy (loser66):

yw